solve by the substitution method.
2x+4y=-2
-8x+y=25
The solution is?

Answer by algebrahouse.com(1659)

2x + 4y = -2
-8x + y = 25------->y = 8x + 25 {added 8x to both sides}

2x + 4(8x + 25) = -2 {substituted 8x + 25, in for y, into top equation}
2x + 32x + 100 = -2 {used distributive property}
34x + 100 = -2 {combined like terms}
34x = -102 {subtracted 100 from both sides}
x = -3 {divided both sides by 34}

y = 8x + 25
y = 8(-3) + 25 {substituted -3, in for x, into y = 8x + 25}
y = -24 + 25 {multiplied 8 by -3}
y = 1 {combined like terms}

x = -3 and y = 1
www.algebrahouse.com